# A comparison between geostatistical and machine learning models for spatio-temporal prediction of PM2.5 data

### v1

@@ 4 Data Analysis
4.1 Feature Selection
Based on the set of features associated with each model and whether the model is considered to be a geostatistical model or non-geostatistical model, we divided all models into four groups. Group 1 contains geostatistical models namely: universal kriging (UK), nearest neighbor Gaussian process (NNGP), and fixed ranked kriging (FRK). Group 2 contains Non-geostatistical methods, namely regression (Reg), random forest (RF), support vector regression (SVR), and ensemble neural network (ENN). The set of features ($p$) for groups 1 and 2 will be the longitude and the latitude for each location. For group 3, we will add the set of the nearest neighbor observations (NNO) for each location and their longitude and latitude as features for Reg, RF, SVR, and ENN. Finally, for group 4 we will add the kriging prediction that we obtained from NNGP to be an additional feature in fitting Reg, RF, SVR, and ENN.
@@ 4.2 Performance Evaluation
To evaluate the performance of each model, we employed five-fold cross-validation. In each fold, PurpleAir (PA) sensors were randomly split spatially to mitigate spatial bias, with 80% for training and 20% for testing each model. There are two types of prediction: estimating PM2.5{2.5} concentrations at locations without PA sensors, which is our focus in this study; and forecasting future PM2.5{2.5} levels at locations with PA sensors. The latter is not the subject of our current investigation. Model performance was assessed using five metrics: root mean squared error (RMSE) in (21), symmetric mean absolute percentage error (SMAPE) in (22), mean absolute deviation (MAD) in (23), the correlation (Cor) between predicted and observed PM2.5concentrations, and 95% prediction coverage. For geostatistical methods, we constructed 95% prediction interval as illustrated in (5).
@@
where $n$ denotes the number of test samples, $Y_{i}$ represents the actual PM2.5{2.5} concentration, and $\hat{Y}{i}$ denotes the corresponding predicted value. These metrics are essential for assessing the accuracy of our predictive models.
@@ 4.3 Results
Table 1 presents the results of comparing geostatistical and non-geostatistical methods as detailed in Section 3. For the initial comparison of prediction performance, we considered geostatistical methods, designated as group 1. Out of the three geostatistical methods, universal kriging demonstrated superior performance across all six metrics, using longitude and latitude as predictors, followed by NNGP and FRK.
In group 2, focusing on non-geostatistical methods, SVR surpassed Reg, RF, and ENN for all metrics, using longitude and latitude as predictors. For group 3, we modified the predictors for the non-geostatistical methods (from group 2) to include the 10 nearest neighbor observations (NNO) of PM2.5 for each location with a one-time lag. Again, SVR outperformed all other methods in this group followed by RF, Reg, and ENN.
In group 4, we adjusted the set of predictors to include predictive values obtained from NNGP (the fastest method from group 1), which significantly enhanced the performance for all machine learning algorithms. SVR maintained the best performance across all metrics.
@@
Figure 5 illustrates how altering the set of predictors can improve RMSE, SMAPE, and Cor across groups 2,3, and 4. It is worth mentioning that, although the RF algorithm comes in second place after SVR, it offers a unique advantage: it allows for the construction of 95% empirical prediction coverage. Through five-fold cross-validation, RF has a predictive coverage of 95% with the addition of Nearest Neighbor Observations NNO and increased to 96% when incorporating the kriging output.
@@
To generate an hourly map of PM2.5 concentrations across California, we constructed a $50\times 50$ grid over 24-hour period. Predictions were made using the top-performing model from each group: Universal Kriging (UK) from group 1, SVR from group 2, SVR with NNO from group 3, and SVR with NNO and Kriging from group 4. Figure 6 displays spatial predictions for PM2.5 concentrations on June 15, 2019 at 2:00 pm. Particularly, panel (d) in 6, associated with the best performing method, reveals that southern California, especially along the coasts, experiences high PM2.5 levels. Stowell et al. (2020) attribute these elevated PM2.5 concentrations in southern California to Santa Ana winds, which can alter temperature, humidity, and wind speed, thereby escalating PM2.5{2.5} levels.
@@ 5 Conclusions and Discussion
This study provides a comprehensive assessment of prevalent geostatistical and machine learning methods for estimating PM2.5 concentrations across spatial and temporal scales, utilizing an extensive network of PurpleAir sensors to enhance prediction accuracy for air quality in California. Among geostatistical methods, universal kriging shows best performance across all six metrics followed by Nearest Neighbor Gaussian Process (NNGP) and Fixed Rank Kriging (FRK). In machine learning methods, Support Vector Regression (SVR) surpassed Regression (Reg), Random Forest (RF), and Ensemble Neural Network (ENN).
@@
While geostatistical methods effectively capture spatiotemporal dependencies, they are constrained by high dimensionality and computational demands. For instance, Universal Kriging (UK) required up to 3600 minutes, compared to 420 minutes for Fixed Rank Kriging (FRK) and only 5 minutes for the Nearest Neighbor Gaussian Process (NNGP). Machine learning algorithms offer rapid implementation and robustness to non-linear relationships, but they generally overlook spatiotemporal dependencies and uncertainty. For the non-geostatistical algorithms, the maximum execution time was 8 minutes recorded for ensemble neural network (ENN).
@@
Our findings highlight the synergistic potential of integrating geostatistical approaches with machine learning to enhance prediction accuracy without relying on assumptions of stationarity, linearity, or normality. Particularly, incorporating predictors from the fastest geostatistical method in our comparison, NNGP, into machine learning models significantly improved prediction quality, reduced RMSE, SMAPE, and MAD, and increased the correlation between observed and predicted PM2.5{2.5} values. Notably, Support Vector Regression (SVR) excelled across all metrics, demonstrating its capability in enhancing the spatiotemporal prediction of PM2.5 concentrations. Overall, our comparison addresses the often-neglected combined spatiotemporal dynamic, which are critical for understanding and managing air quality. More importantly, our findings reveal significant advantages when integrating geostatistical methods with machine learning algorithms.
Earlier research has confirmed that human activities, such as industrial processes, transportation, and agricultural practices significantly impact ground-level PM2.5 concentrations (Bao et al., 2016; Pérez et al., 2010)). Additionally, correlations between land cover/land use (LCLU) types and PM2.5 concentrations suggest that different land uses may reflect varied emission sources (Hoek et al., 2008). Therefore, incorporating meteorological variables, such as temperature, humidity, dew-point temperature, alongside these human factors could further refine predictive models, offering insights into pollution hotspots and assisting policymakers in optimizing air quality monitoring networks.
@@
Looking ahead, future studies could explore the integration of more diverse data sources, including real-time traffic and industrial emissions data, to further contextualize the spatiotemporal dynamics of air pollution. Additionally, advancing algorithmic approaches to more effectively process large-scale environmental data could provide deeper insights into the complex interactions affecting air quality. Such advancements will not only improve predictive accuracy but also enhance our understanding of the environmental and anthropogenic factors driving PM2.5 variability, supporting more targeted and effective air quality management strategies.
@@ Acknowledgments
The authors would like to thank Yawen Guan for helpful discussions and we also appreciate the effort of the Editor, Associate Editor and Referees to improve quality of the manuscript.

### v2

@@ 4 Data Analysis
4.1 Feature Selection and Comparative Framework
Based on the set of features associated with each model and whether the model is considered to be a geostatistical model or non-geostatistical model, we divided all models into four groups. Group 1 contains geostatistical models: universal kriging (UK), nearest neighbor Gaussian process (NNGP), and fixed rank kriging (FRK). Group 2 contains non-geostatistical methods, namely regression (Reg), random forest (RF), support vector regression (SVR), and ensemble neural network (ENN). The set of features ($p$) for groups 1 and 2 will be the longitude and latitude for each location. For group 3, we will add the set of the nearest neighbor observations (NNO) for each location and their longitude and latitude as features for Reg, RF, SVR, and ENN. Finally, for group 4 we will add the kriging that we obtained from NNGP to be an additional feature in the fitting of Reg, RF, SVR, and ENN.
We intentionally limited the predictors to spatial coordinates, nearest neighbor observations, and kriging-based features rather than including additional meteorological or land-use variables. This restriction allows a controlled comparison of model architectures by removing confounding effects of covariate variability. The goal is to isolate the capacity of each modeling approach to capture spatial dependence and short-term variability, rather than to build a comprehensive environmental model. While inclusion of meteorological covariates (e.g., temperature, humidity, wind speed) may further improve accuracy, that direction is beyond the scope of this comparative methodological study.
@@ 4.2 Performance Evaluation
We evaluate the predictive performance across models via five-fold cross-validation by randomly sampling, without replacement, 20% of the PA sensors to be in the test set in each of the five folds. Within each fold, we train each of the models on the data from the remaining 80% of the PA sensors, and we generate predictions at the test sites. This design provides an assessment of short-range predictive accuracy, which is most relevant for the dense PurpleAir network where sensors are geographically close and interpolation is of greater interest than long-range extrapolation. Although spatially blocked validation can reduce spatial dependence between folds, it tends to over-penalize performance when data exhibit fine-scale spatial clustering. Our random spatial splitting thus offers a balanced and fair framework for comparing relative model performance at local scales. We emphasize that the purpose of this validation scheme is to evaluate interpolation ability within the existing sensor network, rather than long-range extrapolation across unsampled regions.
Under this framework, two prediction tasks can be defined. The first is spatial interpolation, estimating PM2.5 concentrations at locations without PA sensors, which is the primary focus of this study. The second is temporal forecasting of PM2.5 at existing sensor sites, which is beyond the scope of the current analysis.
Model performance was assessed using five metrics: root mean squared error (RMSE) in (21), symmetric mean absolute percentage error (SMAPE) in (22), mean absolute deviation (MAD) in (23), correlation (Cor) between predicted and observed PM2.5 concentrations, and 95% prediction coverage. For geostatistical methods, we constructed a 95% prediction interval as illustrated in (5).
@@
where $n$ denotes the number of test samples, $Y_{i}$ represents the actual PM2.5 concentration, and $\hat{Y}{i}$ denotes the corresponding predicted value. These metrics are essential to assess the accuracy of our predictive models.
All model training and evaluation were conducted using the same training–test splits and metrics for consistency. The hyperparameters of the machine learning models were tuned using five-fold cross-validation, while the geostatistical parameters were estimated by maximum likelihood. After cross-validation was used to assess predictive performance, a final version of each model was fit using the full dataset with the same specifications and used for all subsequent analyses. To evaluate model adequacy beyond predictive accuracy, we assessed whether fitted models adequately removed spatial dependence from the residuals. Specifically, residuals at held-out test locations were examined using global Moran’s I statistics, which are standard diagnostics for detecting remaining spatial autocorrelation. All reported performance metrics and residual diagnostics in this manuscript should therefore be interpreted as measures of within-area interpolation performance rather than as assessments of generalization to spatially independent regions.
@@ 4.3 Results
In Group 1 (geostatistical baselines), UK was the strongest (RMSE = 0.3730, SMAPE = 7.802%, MAD = 0.2630, $\mathrm{Cor}=0.8701$) with 95% coverage of 93% but very high computation (3,600 min); NNGP was far faster (5 min) with modestly higher error (RMSE = 0.4076; coverage = 89%), and FRK yielded RMSE = 0.4227 (420 min; coverage = 85%). Non-geostatistical models alone (Group 2) underperformed these baselines (e.g., SVR RMSE = 0.4186, $\mathrm{Cor}=0.6333$). Augmenting machine learning models with nearest-neighbor observations (Group 3) produced large gains at modest cost, for example, SVR+NNO improved to RMSE = 0.1552, SMAPE = 2.740%, MAD = 0.0920, $\mathrm{Cor}=0.9249$ in $\sim$5.11 min. Adding a single NNGP-kriging predictor (Group 4) gave the best overall performance: SVR+NNO+Krig achieved RMSE = 0.0819, SMAPE = 1.330%, MAD = 0.0449, $\mathrm{Cor}=0.9792$ in $\sim$10.93 min; Reg+NNO+Krig and RF+NNO+Krig delivered similar accuracy (RMSE = 0.0826 and 0.0989, respectively) with well-calibrated uncertainty (95% coverage = 97% and 96%). Overall, relative to UK, the top hybrid (SVR+NNO+Krig) reduced RMSE by $\approx$78% (0.3730 to 0.0819) while remaining orders of magnitude faster in computation (about 11 vs. 3,600 min), illustrating a favorable accuracy–compute trade-off for the hybrid approach.
@@
Figure 5 illustrates how altering the set of predictors can improve RMSE, SMAPE, and Cor in groups 2,3 and 4. It is worth mentioning that, although the RF algorithm comes in second place after SVR, it offers a unique advantage: it allows for the construction of 95% empirical prediction coverage. It is important to note that not all models considered in this study provide uncertainty quantification (UQ) in a comparable form. Geostatistical models (UK, FRK, NNGP) yield analytical predictive variances through their Gaussian process formulations, and Random Forest allows empirical uncertainty estimation from the variability among trees. However, Support Vector Regression (SVR) and Ensemble Neural Network (ENN) models are deterministic point predictors that do not natively produce predictive variance. Obtaining UQ for these models would require substantial modification, such as Bayesian SVR, quantile regression, or Bayesian neural networks which falls outside the methodological comparison framework of this paper. For this reason, we report 95% empirical coverage only for models where valid uncertainty estimates are available.
The rightmost column in Table 1 also reports the end-to-end elapsed time for each model, including feature construction (adding nearest-neighbor observations, NNO, for Group 3; and the NNGP kriging feature for Group 4), model fitting, and prediction/evaluation. Relative to Group 2, adding NNO features increased the runtime only modestly, for example, SVR from 1.51 min to 5.11 min and RF from 3.57 min to 6.01 min, while markedly improving accuracy (SVR RMSE from 0.4186 to 0.1552; RF RMSE from 0.4927 to 0.1674). Incorporating the NNGP kriging surrogate in Group 4 produced a further but still modest time increase, for example, SVR+NNO from 5.11 min to 10.93 min, accompanied by additional accuracy gains (SVR RMSE from 0.1552 to 0.0819) and improved calibration for models reporting coverage. Overall, these results indicate a favorable computation-accuracy trade-off: small increases in end-to-end time, fully accounting for upstream feature construction costs, yield substantial predictive benefits. To assess model adequacy and remaining spatial dependence, we evaluated residual spatial autocorrelation for each fitted model using both global Moran’s $I$ statistics. Residuals were computed at held-out test locations as $e_{i}=Y_{i}-\hat{Y}_{i}$. For the geostatistical models that explicitly account for spatial dependence—Universal Kriging (UK), Fixed Rank Kriging (FRK), and NNGP—the residual Moran’s $I$ values were close to zero (UK: I=0.069I=-0.069, FRK: $I=0.021$, NNGP: I=0.084I=-0.084) and statistically non-significant (all $p\geq 0.129$), indicating no detectable residual spatial autocorrelation. For machine-learning models augmented with spatial information through kriging predictions and nearest-neighbor operators, residual spatial dependence was also effectively mitigated. In particular, the hybrid SVR+NNO+Krig, RF+NNO+Krig, Reg+NNO+Krig, and ENN+NNO+Krig models all yielded small Moran’s $I$ values (ranging from 0.082-0.082 to $0.044$) with non-significant $p$-values ($p\geq 0.215$), indicating no detectable residual spatial autocorrelation at the global scale. Consistent with these results, empirical semivariograms of the residuals (shown in the Appendix) do not exhibit a clear increasing trend with distance or a well-defined range, suggesting no detectable large-scale residual spatial dependence at the spatial scales considered. Together, these diagnostics indicate that both the explicitly spatial models and the proposed hybrid learning frameworks effectively mitigate residual spatial dependence. As a result, model performance can be interpreted without confounding effects from unmodeled spatial structure at the spatial scales considered.
@@
To generate an hourly map of PM2.5 concentrations in California, we constructed a $50\times 50$ grid over a 24-hour period. Predictions were made using the highest performing model from each group: UK from group 1, SVR from group 2, SVR with NNO from group 3, and SVR with NNO and Kriging from group 4. Figure 6 displays spatial predictions for PM2.5 concentrations on June 15, 2019 at 2:00 pm. In particular, panel (d) in Figure 6, associated with the best performing method, shows a pronounced hotspot in Southern California. The bright yellow patch near ($34^{\circ}\mathrm{N},117^{\circ}\mathrm{W}$) corresponds to the Riverside–San Bernardino (Inland Empire) region at the eastern edge of the Los Angeles Basin. Elevated PM2.5 in this corridor is consistent with prior literature attributing high levels to basin-confining topography, goods-movement and on-road emissions, episodic inversions/stagnation, and seasonal meteorology (including Santa Ana wind regimes) that modulate temperature, humidity, and wind speed (Stowell et al., 2020).
@@
4.4 Sensitivity Analysis: Validation Strategy
To assess the sensitivity of model comparisons to the choice of validation strategy, we conducted an additional spatially blocked cross-validation analysis for the hybrid Non-Geo+NNO+Krig models. In contrast to random cross-validation, which allows nearby sensors to appear in both training and test sets, spatial cross-validation withholds entire spatial blocks during model fitting and uses them exclusively for evaluation, thereby reducing spatial dependence between folds.
Table 2 compares predictive performance under random and spatial cross-validation. As expected, spatial cross-validation results in higher prediction errors and lower correlations across all models, reflecting the increased difficulty of predicting into spatially independent regions rather than interpolating among nearby sensors. Importantly, however, the relative performance ordering of the models remains consistent across validation strategies, with hybrid approaches combining non-spatial predictors, nearest-neighbor operators, and kriging continuing to outperform alternative methods.
These results indicate that while random cross-validation provides optimistic estimates of absolute performance, the qualitative conclusions regarding comparative model performance are robust to the choice of validation strategy.
Random CV Spatial CV
Method RMSE SMAPE MAD Cor 95% Cov RMSE SMAPE MAD Cor 95% Cov
NNGP 0.4076 9.117% 0.3174 0.7925 89% 0.53 11.0% 0.380 0.69 92%
Reg+NNO+Krig 0.0826 1.390% 0.04697 0.9781 97% 0.18 3.5% 0.120 0.92 85%
RF+NNO+Krig 0.0989 1.458% 0.0493 0.9694 96% 0.14 2.4% 0.075 0.90 94%
SVR+NNO+Krig 0.0819 1.330% 0.0449 0.9792 0.12 2.1% 0.065 0.92
ENN+NNO+Krig 0.0994 2.149% 0.0552 0.9690 0.22 4.5% 0.150 0.89
Table 2: Comparison of hybrid Non-Geo+NNO+Krig models under random cross-validation (Random CV) and spatial cross-validation (Spatial CV).
@@ 5 Conclusions and Discussion
This study provides a comprehensive assessment of prevalent geostatistical and machine learning methods for estimating PM2.5 concentrations across spatial and temporal scales, utilizing an extensive network of PurpleAir sensors to enhance prediction accuracy for air quality in California.
@@
While geostatistical methods effectively capture spatiotemporal dependencies, they are constrained by high dimensionality and computational demands. For instance, Universal Kriging required up to 3600 minutes, compared to 420 minutes for Fixed Rank Kriging and only 5 minutes for the Nearest Neighbor Gaussian Process. For the non-geostatistical machine learning algorithms, the maximum execution time was about 6 minutes recorded for ensemble neural network. Machine learning algorithms offer rapid implementation and robustness to non-linear relationships, but they generally overlook spatiotemporal dependencies and uncertainty.
@@
Our findings highlight the synergistic potential of integrating geostatistical approaches with machine learning to enhance prediction accuracy without relying on assumptions of stationarity, linearity, or normality. Particularly, incorporating predictors from the fastest geostatistical method in our comparison, Nearest Neighbor Gaussian Process, into machine learning models significantly improved prediction quality, reduced RMSE, SMAPE, and MAD, and increased the correlation between observed and predicted PM2.5 values. Among the hybrid models, SVR+NNO+Krig achieved the lowest point-prediction error across evaluation metrics. However, models such as Reg+NNO+Krig and RF+NNO+Krig provided comparable predictive accuracy while also delivering calibrated uncertainty estimates, which may be preferable in operational or decision-support contexts. Overall, our comparison addresses the often-neglected combined spatiotemporal dynamics, which are critical for understanding and managing air quality. More importantly, our findings reveal significant advantages when integrating geostatistical methods with machine learning algorithms.
Despite these strengths, several limitations should be acknowledged. First, we used the Barkjohn et al. (2020) calibration formula for correcting PurpleAir data, which, although widely adopted, may not fully capture spatial or humidity-driven variation in sensor bias. More recent dynamic or site-specific calibration models could further improve data quality, particularly during wildfire events or high-humidity periods. Second, while our primary evaluation focused on short-range interpolation accuracy using random spatial cross-validation within a dense sensor network, we additionally conducted spatially blocked cross-validation as a sensitivity analysis. As expected, spatial blocking produced higher error metrics and lower correlations, reflecting the increased difficulty of predicting into spatially independent regions rather than interpolating among nearby sensors. However, the relative performance ordering of the models was preserved, indicating that the comparative advantages of the hybrid framework are robust to the choice of validation strategy. Nevertheless, absolute performance under spatial generalization remains more conservative than under interpolation settings, and applications targeting geographically distinct regions may require validation schemes tailored to broader extrapolation contexts. Third, while deterministic machine learning models such as Support Vector Regression and Ensemble Neural Network demonstrated excellent point predictive accuracy, they lack intrinsic uncertainty quantification. Future extensions could explore probabilistic machine learning formulations, such as Bayesian support vector regression or Monte Carlo dropout neural networks, to provide calibrated predictive uncertainty alongside point predictions and improve interpretability.
Incorporating additional environmental and anthropogenic covariates, such as meteorological variables, land-use characteristics, and emissions-related indicators, could further enhance predictive performance and improve the identification of pollution hotspots (Bao et al., 2016; Pérez et al., 2010; Hoek et al., 2008). Relative to recent advances in spatiotemporal deep learning and data fusion approaches that integrate satellite retrievals, chemical transport models, and sensor data, our hybrid geostatistical machine learning framework offers a complementary perspective. While large-scale data fusion methods often prioritize broad spatial generalization, our framework emphasizes superior local interpolation performance within dense sensor networks and, for models providing predictive variance (e.g., geostatistical and RF-based approaches), interpretable spatial uncertainty quantification.
@@
Looking ahead, future studies could explore the integration of more diverse data sources, including real-time traffic and industrial emissions data, to further contextualize the spatiotemporal dynamics of air pollution. Additionally, advancing algorithmic approaches to more effectively process large-scale environmental data could provide deeper insights into the complex interactions affecting air quality. Such advancements will not only improve predictive accuracy but also enhance our understanding of the environmental and anthropogenic factors driving PM2.5 variability, supporting more targeted and effective air quality management strategies. When combined with probabilistic modeling frameworks that provide uncertainty quantification, the integration of richer environmental covariates and large-scale data streams would further enhance model interpretability and decision relevance for air-quality management.
@@ Acknowledgments
The authors would like to thank Tamer Oraby, Yawen Guan for helpful discussions. We also appreciate the effort of the Editor, Associate Editor and Referees to improve quality of the manuscript.
Appendix A Appendix
(a) UK
(b) NNGP
(c) FRK
(d) Reg+NNO+Krig
(e) RF+NNO+Krig
(f) SVR+NNO+Krig
Figure 7: Residual semivariogram plots for (a) UK, (b) NNGP, (c) FRK, (d) Reg+NNO+Krig, (e) RF+NNO+Krig, and (f) SVR+NNO+Krig.